If a football is kicked straight up with an initial velocity of 64 ft/sec from a height of 5 ft, then its height above the earth is a function of time given by h(t)=-16t^2+64t+5.
What is the maximum height reached by the ball in ft?

Answer by Alan3354(69443)

If a football is kicked straight up with an initial velocity of 64 ft/sec from a height of 5 ft, then its height above the earth is a function of time given by h(t)=-16t^2+64t+5.
What is the maximum height reached by the ball in ft?
========================================
It's the vertex of the parabola h(t)=-16t^2+64t+5.
---------------
The time to max height is t = -b/2a = -64/(-32) = 2 seconds
h(2) = -16*4 + 64*2 + 5
= 69 feet